The measure of the supplement of an angle is 30 less than twice the measure of the angle.Find the measure of each.
Answer by mananth(16949)

let angle be x
the complement is 2x-30
..
x+2x-30=90
3x-30=90
3x=90+30
3x=120
x= 40.
the other angle will be 50 deg.
..